### v2

# 22-quasi-perfect Lee codes and abelian Ramanujan graphs: a new construction and relationship

Abstract.
This paper presents a new explicit infinite family of 2-quasi-perfect $p$-ary Lee codes of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ for $q=p^{k}\geq 14$, $p\geq 5$ a prime. Our codes are derived from the generating set $H_{q}=\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$ of the finite field $\mathbb{F}_{q^{2}}$. Furthermore, we bridge between 2-quasi-perfect Lee codes constructed by Mesnager, Tang, and Qi and well-known abelian Ramanujan graphs, specifically Li’s graphs and finite Euclidean graphs, providing a unified theoretical framework for these families.
Key words and phrases:
Quasi-perfect Lee codes, abelian Ramanujan graphs, Li’s graphs, Finite Euclidean graphs.
2020 Mathematics Subject Classification:
94B05, 94B27, 05C48
1. Introduction
A Lee code is an error-correcting code under the Lee distance. While finding perfect Lee codes is a classical problem, from a broader theoretical perspective, the focus has shifted to the construction of “good” codes, such as quasi-perfect Lee codes. Constructing linear quasi-perfect Lee codes over finite fields represents a fascinating intersection of coding theory, number theory, algebraic geometry, and combinatorics. In this context, Camarero and Martínez [3] constructed $2$-quasi-perfect $p$-ary Lee codes of length $2\lfloor\frac{p}{4}\rfloor$ for arbitrary primes $p\equiv\pm 5\pmod{12}$, which proves the existence of quasi-perfect Lee codes for arbitrarily large length. For an odd prime $p$, Mesnager, Tang, and Qi [15] formulated an equivalent condition for the existence of $2$-quasi-perfect $p$-ary linear codes of length $n$, and provided several explicit families of $2$-quasi-perfect Lee codes by generalizing the Camarero-Martínez construction and analyzing the solution sets of specific polynomial equations. Moreover, the Camarero-Martínez and Mesnager-Tang-Qi constructions also produce Cayley graphs over the additive groups of finite fields that are abelian Ramanujan graphs. It is worth mentioning that these graphs are optimal pseudorandom graphs in a sense of spectral gap, where their short diameters crucially affect the covering property of the corresponding Lee codes.
Nevertheless, to the best of our knowledge, constructing quasi-perfect Lee codes of arbitrarily large length remains a challenging open problem. A unified theoretical framework that rationalizes why certain algebraic structures yield such codes is notably absent; consequently, existing constructions often appear as isolated families derived from specific polynomials without clear intuitive grounding. Furthermore, while the small diameter of abelian Ramanujan graphs inherently favors the covering radius of associated Lee codes, achieving perfect or quasi-perfect status necessitates stricter structural constraints. Determining precisely which abelian Ramanujan graphs yield these optimal codes remains a significant open question.
In this paper, we first exhibit a new and simple construction for $2$-quasi-perfect linear Lee codes of arbitrarily large length. Specifically, we investigate the $p$-ary linear code derived from the set $H_{q}\coloneq\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$, where $p\geq 5$ is a prime and $q=p^{k}$, and prove that this code is a $2$-quasi-perfect code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ provided that $q\geq 14$ (Theorem 4). It is worth mentioning that this construction comes from a recent construction of abelian almost Ramanujan graphs by Forey, Fresán, Kowalski and Wigderson [7], and the structure of the codes and associated Cayley graphs are different from known ones. Moreover, we bridge the Mesnager-Qi-Tang construction of Lee codes to well-studied abelian Ramanujan graphs in combinatorics and number theory, namely, Li’ s graphs introduced by Winnie Li ([11]), and finite Euclidean graphs ([14]). This provides a unified description of the families of Lee codes by the Mesnager-Qi-Tang construction (Propositions 16 and 18).
The remainder of the paper is organized as follows. Section 2 reviews preliminaries on finite fields, Lee codes, Cayley and Ramanujan graphs. Section 4 details our construction and the proof of the $2$-quasi-perfectness employing algebraic geometry. Section 5 discusses a connection between the Mesnager-Qi-Tang codes and abelian Ramanujan graphs, namely, Li’s graphs and finite Euclidean graphs. We conclude this paper with several remarks in Section 6.
2. Preliminaries
2.1. Finite fields
Let $p$ be a prime, and $q=p^{k}$, where $k\geq 1$ is an integer. Throughout this paper, $\mathbb{F}_{q}$ denotes the finite field of order $q$. Note that $\mathbb{F}_{p}$ is the residue ring $\mathbb{Z}/(p\mathbb{Z})$, and it is a subfield of $\mathbb{F}_{q}$. The additive group of $\mathbb{F}_{q}$, denoted by $\mathbb{F}_{q}^{+}$, has the structure of the $k$-dimensional vector space over $\mathbb{F}_{p}$, and accordingly, for each integer $n\geq 1$, $\mathbb{F}_{q}^{n}\coloneq\{(a_{1},a_{2},\ldots,a_{n})\mid a_{i}\in\mathbb{F}_{q},i=1,2,\ldots,n\}$ is a $kn$-dimensional vector space over the same field $\mathbb{F}_{p}$. In addition, let $\mathbb{F}_{q}^{*}\coloneq\mathbb{F}_{q}\setminus\{0\}$, which is the multiplicative group of $\mathbb{F}_{q}$. For $\ell\geq 1$, the field norm ${\rm N}_{q^{\ell}/q}$ from $\mathbb{F}_{q^{\ell}}^{*}$ to $\mathbb{F}_{q}^{*}$ is a surjective homomorphism defined as ${\rm N}_{q^{\ell}/q}(z)\coloneq z^{1+q+\cdots+q^{\ell-1}}$. Also, for $m\geq 1$, let $GL(m,\mathbb{F}_{q})$ denote the general linear group of rank $m$ over $\mathbb{F}_{q}$, consisting of all invertible $m\times m$ matrices over $\mathbb{F}_{q}$. For further details of finite fields, see, e.g., [12].
2.2. Lee codes
Let $p\geq 3$ be a prime. A $p$-ary (linear) Lee code $\mathcal{C}$ of length $n$ is a linear subspace of the vector space $\mathbb{F}_{p}^{n}$, where the error correction is formulated via the following Lee distance: for any $\mathbf{a},\mathbf{b}\in\mathbb{F}_{p}^{n}$,
$d_{L}(\mathbf{a},\mathbf{b})\coloneq\sum_{i=1}^{n}\min\{|s|\mid s\equiv a_{i}-b_{i}\pmod{p},\>s\in\mathbb{Z}\}.$
An element of a linear code $\mathcal{C}$ is called a codeword of $\mathcal{C}$. A code $\mathcal{C}$ is said to be $t$-error-correcting if $t$ is the largest integer such that for any $\mathbf{w}\in\mathbb{F}_{p}^{n}$, there exists at most one codeword $\mathbf{c}\in\mathcal{C}$ with $d_{L}(\mathbf{w},\mathbf{c})\leq t$, where such $t$ is called the packing radius of $\mathcal{C}$. A code $\mathcal{C}$ is said to be $r$-covering if $r$ is the smallest integer such that for any $\mathbf{w}\in\mathbb{F}_{p}^{n}$, there exists at least one codeword $\mathbf{c}\in\mathcal{C}$ with $d_{L}(\mathbf{w},\mathbf{c})\leq r$, where such $r$ is called the covering radius of $\mathcal{C}$ We say a code $\mathcal{C}$ is $t$-perfect if it is both $t$-error-correcting and $t$-covering. A famous Golomb-Welch conjecture states that a $t$-perfect code of length $n$ exists only if $t=1$ or $n=1,2$. Instead, the best alternative is known as a $t$-quasi-perfect code, where a code $\mathcal{C}$ is said to be $t$-quasi-perfect if it is both $t$-error-correcting and $(t+1)$-covering. Even for $t$-quasi-perfect codes with $t\geq 2$, there are few explicit constructions, and it is quite challenging to obtain explicit constructions for $t$-quasi-perfect codes.
2.3. Cayley graphs and Ramanujan graphs
For a finite group $G$ with identity $1$, and an inverse-closed subset $S\subset G\setminus\{1\}$ (which means that $s^{-1}\in S$ for any $s\in S$), the Cayley graph $Cay(G,S)$ on $G$ with respect to $S$ is the graph with vertex set $G$ in which two vertices $g$ and $h$ are adjacent if and only if $gh^{-1}\in S$. Note that $Cay(G,S)$ is an undirected $|S|$-regular graph.
The adjacency matrix $A_{X}$ of a simple $N$-vertex graph $X$ is an $N\times N$ $\{0,1\}$-matrix where the $(u,v)$-entry is $1$ if and only if $u$ and $v$ are adjacent in $X$, and hence, the matrix is symmetric, implying that all eigenvalues are real. By the Perron-Frobenius theorem, for any $d$-regular graph $X$, all eigenvalues of $A_{X}$ are in the interval [d,d][-d,d], where the largest eigenvalue is $d$, and d-d is an eigenvalue if and only if $X$ is bipartite. We call $\pm d$ trivial eigenvalues, and set $\lambda(X)$ as the largest absolute value of non-trivial eigenvalues of $A_{X}$. A $d$-regular NN-vertex graph $X$ is called a Ramanujan graph if $\lambda(X)\leq 2\sqrt{d-1}$. Similarly, we say $X$ is an almost Ramanujan graph if $\lambda(X)\leq(2+o(1))\sqrt{d-1}$ as $N\to\infty$ If $d$ is fixed, then it is well-known that $\liminf_{N\to\infty}\lambda(X)\geq 2\sqrt{d-1}$ by the Alon-Boppana bound (e.g. [5]), and even if $d$ grows as $N\to\infty$ (say, $\omega(1)=d<\frac{N}{2}$), one can prove that $\lambda(X)=\Omega(\sqrt{d})$ as $N\to\infty$. Hence, Ramanujan graphs realize the optimally small non-trivial eigenvalues in absolute value (up to constant). It is known that (almost) Ramanujan graphs enjoy very rich properties such as high global connectivity, pseudo-randomness, and rapid mixing. An abelian Ramanujan graph is a Cayley graph over an abelian group which is a Ramanujan graph. For a more detailed exposition of Cayley and Ramanujan graphs, see, for instance, [5].
3. A generic construction
Recall a generic construction for $2$-quasi-perfect $p$-ary Lee codes by Mesnager, Tang and Qi in [15].
Definition 1 ([15]).
Let $\Gamma$ be a vector space over $\mathbb{F}_{p}$ which is spanned by a set $H=\{\pm\beta_{1},\pm\beta_{2},\ldots,\pm\beta_{n}\}$ of non-zero vectors. Then define the $p$-ary code $\mathcal{C}(\Gamma,H)$ as
$\mathcal{C}(\Gamma,H)\coloneq\{(c_{1},\ldots,c_{n})\in\mathbb{F}_{p}^{n}\mid c_{1}\beta_{1}+\cdots+c_{n}\beta_{n}=\mathbf{0}\}.$
For subsets $A,B$ of a vector space, let $A\coloneq\{-a\mid a\in A\}$, $A+B\coloneq\{a+b\mid a\in A,b\in B\}$. For an integer $i\geq 0$, define
$A^{(i)}\coloneq\begin{cases}\{0\}&\text{if $i=0$,}\\ A&\text{if $i=1$,}\\ A^{(i-1)}+A&\text{if $i\geq 2$.}\end{cases}$
Lemma 2 (Proposition 6 in [15]).
Let $p\geq 5$ be an odd prime. For $\Gamma$ and $H$ satisfying the assumptions of Definition 1, we have the following.
  1. $(1)$
    $\mathcal{C}(\Gamma,H)$ is a $2$-perfect $p$-ary Lee code of length $n$ and dimension $n-\dim_{\mathbb{F}_{p}}(\Gamma)$ if and only if $(\bigcup_{i=0}^{2}H^{(i)})=2n^{2}+2n+1=\#\Gamma$.
  2. $(2)$
    $\mathcal{C}(\Gamma,H)$ is a $2$-quasi-perfect $p$-ary Lee code of length $n$ and dimension $n-\dim_{\mathbb{F}_{p}}(\Gamma)$ if and only if all of the following hold:
    • $(\bigcup_{i=0}^{2}H^{(i)})=2n^{2}+2n+1$,
    • $\bigcup_{i=0}^{3}H^{(i)}=\Gamma$,
    • $2n^{2}+2n+1<\#\Gamma<\frac{1}{3}(2n+1)(2n^{2}+2n+3)$.
Remark 3.
Consider the Cayley graph $Cay(\Gamma,H)$, where $\Gamma$ is also an abelian group. Then $H^{(i)}$ is the set of all vertices to which the distance from the zero-vector $\mathbf{0}$ is $i$. Hence, the second condition of Lemma 2 (2) shows that the diameter of $Cay(\Gamma,H)$ is $3$. On the other hand, it is known (e.g. [4]) that for any $d$-regular NN-vertex graph $X$, its diameter is at most $\lceil\frac{\log(N-1)}{\log(d/\lambda(X))}\rceil$, and hence, among all $d$-regular NN-vertex graphs, $d$-regular NN-vertex Ramanujan graphs minimize the upper bound for the diameter. Thus, for constructing a $2$-quasi-perfect code $\mathcal{C}(\Gamma,H)$, it is hopeful to choose $H$ so that $Cay(\Gamma,H)$ is an abelian Ramanujan graph. On the other hand, the conditions of Lemma 2 (2) strictly restrict the structure of $Cay(\Gamma,H)$, and hence finding suitable $H$ would be quite difficult in general.
4. A new explicit construction of $2$-quasi-perfect Lee codes
For a prime $p\geq 5$ and a power $q=p^{k}$ with $k\geq 1$, let $H_{q}\coloneq\{(a,a^{3})\mid a\in\mathbb{F}_{q}^{*}\}$. Notice that as $q$ is odd, $H_{q}$ is even and $H_{q}=2\cdot\frac{q-1}{2}$. In addition, $H_{q}$ is the set of points on the cubic curve $y=x^{3}$.
Here is a new explicit $2$-quasi-perfect Lee codes.
Theorem 4.
Let $p\geq 5$ be a prime. Suppose that a power $q=p^{k}$ with $k\geq 1$ satisfies $q\geq 14$. Then, the code $\mathcal{C}(\mathbb{F}_{q}^{2},H_{q})$ is a $2$-quasi-perfect $p$-ary Lee code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$.
Remark 5.
By the Weil sum estimation (e.g. [12]), $Cay(\mathbb{F}_{q^{2}}^{+},H_{q})$ is a (q1)(q-1)-regular almost Ramanujan graph. This graph was found by Forey, Fresán, Kowalski and Wigderson in [7], and not isomorphic to the abelian Cayley graph associated with known 22-quasi-perfect Lee codes [2, 15].
Remark 6.
As shown in Theorem 14 (2), Mesnager, Qi and Tang provided a $2$-quasi-perfect $p$-ary Lee code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$ for primes $p\equiv 5,11\pmod{12}$ and $q=p^{k}>12$. For these $p$ and $q$, the code in Theorem 4 is not equivalent to their code although both have the same parameters. Moreover, our codes exhibit new parameters for primes $p\equiv 1,7\pmod{12}$.
To illustrate the distinct algebraic structures of our new codes, we exhibit their parity-check matrices alongside those of the Mesnager-Tang-Qi codes for several small parameters.
Example 7.
For $p=17\equiv 5\pmod{12}$ and $k=1$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{17}^{2},H_{17})$ is
$\begin{pmatrix}1&2&3&4&5&6&7&8\\ 1&8&10&13&6&12&3&2\end{pmatrix}$
while the parity check matrix of the Mesnager-Qi-Tang code in Theorem 14 (2) is
$\begin{pmatrix}1&2&3&4&5&6&7&8\\ 1&9&6&13&7&3&5&15\end{pmatrix}$.
Example 8.
For $p=19\equiv 7\pmod{12}$ and $k=1$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{19}^{2},H_{19})$ is
$\begin{pmatrix}1&2&3&4&5&6&7&8&9\\ 1&8&8&7&11&7&1&18&7\end{pmatrix}$.
Example 9.
For $p=23\equiv 11\pmod{12}$ and $k=1$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{23}^{2},H_{23})$ is
[…]
while the parity check matrix of the Mesnager-Qi-Tang code in Theorem 14 (2) is
[…]
Example 10.
For $p=5$ and $k=2$, the parity check matrix of the code $\mathcal{C}(\mathbb{F}_{5}^{4},H_{5^{2}})$ is
$\left(\begin{array}[]{rrrrrrrrrrrr}1&3&1&0&3&3&2&1&2&0&1&1\\ 0&4&4&1&3&1&0&3&3&2&1&2\\ 1&0&2&0&4&0&3&0&1&0&2&0\\ 0&1&0&2&0&4&0&3&0&1&0&2\end{array}\right).$
Here, we express $\mathbb{F}_{25}$ as $\mathbb{F}_{5}[\sqrt{2}]$ by taking the irreducible polynomial $x^{2}-2$, and choose $3+4\sqrt{2}$ as a primitive element of $\mathbb{F}_{25}$.
The proof of Theorem 4 is based on careful verifications for the equivalent condition in Lemma 2 (2), which are completed by the following series of technical lemmas.
Lemma 11.
For any prime $p\geq 5$ and power $q=p^{k}$ with $k\geq 1$, we have
$\#\bigcup_{i=0}^{2}H_{q}^{(i)}=2\cdot\Bigl(\frac{q-1}{2}\Bigr)^{2}+2\cdot\Big(\frac{q-1}{2}\Big)+1$
Proof.
First we show that $H_{q}^{(2)}=\frac{(q-1)^{2}}{2}+1$. It suffices to count $(a,b)\in\mathbb{F}_{q}^{2}$ such that there are $x,y\in\mathbb{F}_{q}^{*}$ with
(4.1) $\begin{cases}a=x+y\\ b=x^{3}+y^{3}\end{cases}$
Note that $a=0$ if and only if y=xy=-x, equivalent to $b=0$, implying that there are exactly $\frac{q-1}{2}$ unordered pair $\{x,y\}$ satisfying (4.1) with $(a,b)=(0,0)$. If $a\neq 0$, then $(a,b)$ corresponds to an unordered pair $\{x,y\}$ satisfying (4.1) and $x+y\neq 0$ by a one-to-one mapping. Indeed, by (4.1), we have $xy=\frac{a^{3}-b}{3a}$, and hence $x,y$ must be roots of the quadratic equation $t^{2}-at+\frac{a^{3}-b}{3a}=0$. The total number of unordered pairs $\{x,y\}$ from $\mathbb{F}_{q}^{*}$ (allowing $x=y$) is equivalent to choosing 2 elements from q1q-1 elements with replacement, which is $\binom{(q-1)+2-1}{2}=\binom{q}{2}=\frac{q(q-1)}{2}$. Thus, the number of unordered pairs with $x+y\neq 0$ is:
$\frac{q(q-1)}{2}-\frac{q-1}{2}=\frac{(q-1)^{2}}{2}.$
Hence, we conclude that $H_{q}^{(2)}=\frac{(q-1)^{2}}{2}+1$ by adding $(0,0)$. The proof completes after showing that $H_{q}\cap H_{q}^{(2)}=\emptyset$ as $H_{q}^{(0)}\subset H_{q}^{(2)}$ and $H_{q}^{(1)}=\#H_{q}=q-1$. Suppose that there exists $(a,a^{3})\in H_{q}$ such that there are $x,y\in\mathbb{F}_{q}^{*}$ with
$\begin{cases}a=x+y\\ a^{3}=x^{3}+y^{3}\end{cases}$
This implies that $3axy=0$, which is impossible as $x,y,a\neq 0$ and $p\geq 5$. ∎
Lemma 12.
Let $p\geq 5$ be a prime. Suppose that a power $q=p^{k}$ with $k\geq 1$. Then the following polynomial $F_{a,b}(x,y)$ over $\mathbb{F}_{q}$ is absolutely irreducible unless $b=a^{3}$.
$F_{a,b}(x,y)\coloneq xy(x+y)$ $-a(x+y)^{2}+a^{2}(x+y)+\frac{b-a^{3}}{3}$
If $b=a^{3}$, then for any $a\in\mathbb{F}_{q}$, we have
$F_{a,a^{3}}(x,y)=(x+y)(x-a)(y-a).$
Proof.
As the statement for the case $b=a^{3}$ is trivial, suppose $b\neq a^{3}$ and $F_{a,b}$ is not absolutely irreducible. Then, $F_{a,b}$ must have a linear factor which is either x+ycx+y-c, xcx-c or ycy-c for some constant $c$. If x+ycx+y-c is a factor of $F_{a,b}$, then substituting $x+y=c$ into $F_{a,b}(x,y)=0$ yields that $cxy-ac^{2}+a^{2}c+\frac{b-a^{3}}{3}=0$ holds for any $x$ and $y$. Hence we have $b=a^{3}$, a contradiction. If xcx-c is a factor of $F_{a,b}$, then it holds for any $y$ that $cy(c+y)-a(c+y)^{2}+a^{2}(c+y)+\frac{b-a^{3}}{3}=0$. Hence we have $c=a$ and $b=a^{3}$, which cannot occur by the assumption. By the same discussion, ycy-c cannot be a factor of $F_{a,b}$, completing the proof. ∎
Lemma 13.
Let $p\geq 5$ be a prime. Suppose that a power $q=p^{k}$ with $k\geq 1$ satisfies $q\geq 14$. Then we have
$\bigcup_{i=0}^{3}H_{q}^{(i)}=\Gamma$
Proof.
As $\{(0,0)\}=H_{q}^{(0)}$, it suffices to prove that for any $(a,b)\in\mathbb{F}_{q}^{2}\setminus\{(0,0)\}$, there are $x,y,z\in\mathbb{F}_{q}^{*}$ such that
$\begin{cases}a=x+y+z\\ b=x^{3}+y^{3}+z^{3}\end{cases}$
Then, by substituting z=axyz=a-x-y into the second equation, we have the cubic curve $C_{a,b}:F_{a,b}(x,y)=0$ of genus one. Hence, desired elements $x,y,z\in\mathbb{F}_{q}^{*}$ exist if there exists a point $(x,y)$ on the curve $C_{a,b}$ with $x,y\neq 0$ and $x+y\neq a$.
If $b\neq a^{3}$, then $F_{a,b}$ is absolutely irreducible by Lemma 12. In addition, on the projective space $\mathbb{P}^{2}_{\mathbb{F}_{q}}$ of dimension $2$, the projective points $(X:Y:Z)$ with $Z=0$ are $(1:0:0),(0:1:0)$ and (1:1:0)(1:-1:0) which must be excluded. Thus, by the Hasse-Weil bound (e.g. [17]), the number of points of $C_{a,b}$ is at least $q-2\sqrt{q}-2$. Note that $C_{a,b}$ cannot intersect with the line $x+y=a$ if $b\neq a^{3}$ by the proof of Lemma 12, and each of the lines $x=0$ and $y=0$ intersects with $C_{a,b}$ at no more than two points. Hence, if $b\neq a^{3}$, there exists a point $(x,y)$ with $x,y\neq 0$ and $x+y\neq a$ if $(q-2\sqrt{q}-2)-2\cdot 2>0$, which holds as $q\geq 14$.
If $b=a^{3}$, as shown in Lemma 12, the curve $C_{a,b}$ contains the line $x+y=0$. Since $a\neq 0$ holds by the assumption, all points (t,t)(t,-t), $t\in\mathbb{F}_{q}^{*}$, on $C_{a,b}$ are desired ones.
These complete the proof. ∎
Proof of Theorem 4.
Notice that $\mathbb{F}_{q}^{2}$ is a vector space over $\mathbb{F}_{p}$ with $\dim_{\mathbb{F}_{p}}(\mathbb{F}_{q}^{2})=2k$. Also, $H_{q}$ clearly spans $\mathbb{F}_{q}^{2}$ by Lemma 13. By taking $n=\frac{q-1}{2}$, all the conditions of Lemma 1 are satisfied by Lemmas 11 and 13, completing the proof. ∎
5. The Mesnager-Tang-Qi codes and abelian Ramanujan graphs
Now recall the explicit $2$-quasi-perfect Lee codes by Mesnager, Tang and Qi in [15]. For an odd prime $p$, let $q=p^{k}$ where $k$ is an integer. Then, the field extension $\mathbb{F}_{q^{2}}$ of $\mathbb{F}_{q}$ is expressed as $\mathbb{F}_{q^{2}}=\mathbb{F}_{q}[\sqrt{\delta}]=\{x+y\sqrt{\delta}\mid x,y\in\mathbb{F}_{q}\}$ where $\delta\in\mathbb{F}_{q}^{*}$ is a non-square. For $z=x+y\sqrt{\delta}\in\mathbb{F}_{q^{2}}$, let $\overline{z}\coloneq x-y\sqrt{\delta}$ and $\mathcal{N}(z)\coloneq z\cdot\overline{z}=x^{2}-\delta y^{2}$. Then define
$H_{+}\coloneq\{z\in\mathbb{F}_{q^{2}}\mid\mathcal{N}(x+y\sqrt{\delta})=1\},$
$H_{-}\coloneq\{(x,y)\in\mathbb{F}_{q}^{2}\mid xy=1\}.$
Clearly, both $H_{+}$ and $H_{-}$ are subsets of $\mathbb{F}_{q}^{2}$.
Theorem 14 (Theorem 8 in [15]).
For an odd prime $p$, let $q=p^{k}$ where $k$ is an integer. Then we have the following.
  1. $(1)$
    If either $p\equiv 1,7\pmod{12}$, or $p\geq 5$ and $k$ is even, then the code $\mathcal{C}(\mathbb{F}_{q}^{2},H_{+})$ is a $2$-quasi-perfect $p$-ary code of length $\frac{q+1}{2}$ and dimension $\frac{q+1}{2}-2k$.
  2. $(2)$
    If $p\equiv 5,11\pmod{12}$ and $q>12$, then the code $\mathcal{C}(\mathbb{F}_{q}^{2},H_{-})$ is a $2$-quasi-perfect $p$-ary code of length $\frac{q-1}{2}$ and dimension $\frac{q-1}{2}-2k$.
The first observation here is that the code in Theorem 14 (1) is in fact obtained by the generating set of the following abelian Ramanujan graph, called Li’s graph.
Definition 15 ([11]).
For any odd prime power $q$, let $N_{2}:=\{z\in\mathbb{F}_{q^{2}}\mid{\rm N}_{q^{2}/q}(z)=1\}$. Then, Li’s graph over $\mathbb{F}_{q^{2}}$ is the Cayley graph $Cay(\mathbb{F}_{q^{2}}^{+},N_{2})$.
It is known in graph theory and number theory that Li’s graphs are Ramanujan graphs by Deligne’s bound for Kloosterman sums over $\mathbb{F}_{q^{2}}$ ([11, 6]).
As observed in [2] for $q=p\equiv 3\pmod{4}$, one can easily verify:
Proposition 16.
For any odd prime power $q$, it holds that $N_{2}=H_{+}$. Hence, the code $\mathcal{C}(\mathbb{F}_{q}^{2},H_{+})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},N_{2})$.
Proof.
It suffices to prove that $\mathcal{N}$ coincides with ${\rm N}_{q^{2}/q}$. For any $z=x+y\sqrt{\delta}\in\mathbb{F}_{q^{2}}$, we have
${\rm N}_{q^{2}/q}(z)$ $=(x+y\sqrt{\delta})(x+y\sqrt{\delta})^{q}$
$=(x+y\sqrt{\delta})(x^{q}+y^{q}(\sqrt{\delta})^{q})$
$=(x+y\sqrt{\delta})(x+y(\sqrt{\delta})^{q})$
$=(x+y\sqrt{\delta})(x+y(-\sqrt{\delta}))$
$=z\cdot\overline{z}=\mathcal{N}(z),$
where $(\sqrt{\delta})^{q}=-\sqrt{\delta}$ as $\delta^{q-1}=1$ and $\delta^{\frac{q-1}{2}}\neq 1$ as $\delta$ is non-square. ∎
The second observation is that all codes in Theorem 14 are unified by the concept of finite Euclidean graphs.
Definition 17 ([1, 14]).
For any odd prime power $q$, let $Q(x,y)$ be a non-degenerate quadratic form over $\mathbb{F}_{q}$, and define $E_{Q}\coloneq\{(x,y)\in\mathbb{F}_{q}^{2}\mid Q(x,y)=1\}$. The unit finite Euclidean graph associated with $Q$ is the Cayley graph $\mathrm{Cay}(\mathbb{F}_{q^{2}}^{+},E_{Q})$.
Let $Q_{+}(x,y)\coloneq x^{2}-\delta y^{2}$ and $Q_{-}(x,y)\coloneq xy$, where $\delta\in\mathbb{F}_{q}^{*}$ is a non-square element. Note that $Q_{+}$ and $Q_{-}$ are inequivalent, and any non-degenerate quadratic form $Q$ over $\mathbb{F}_{q}$ is equivalent to either $Q_{+}$ or $Q_{-}$. Recall that two quadratic forms $Q_{1}(\mathbf{x})$ and $Q_{2}(\mathbf{x})$ are equivalent if there exists $A\in\mathrm{GL}(2,\mathbb{F}_{q})$ such that $Q_{2}(\mathbf{x})=Q_{1}(A\mathbf{x})$ for all $\mathbf{x}\in\mathbb{F}_{q}^{2}$. Consequently, the finite Euclidean graph $\mathrm{Cay}(\mathbb{F}_{q^{2}}^{+},E_{Q})$ is isomorphic to either $G_{+}\coloneq\mathrm{Cay}(\mathbb{F}_{q^{2}}^{+},E_{Q_{+}})$ or $G_{-}\coloneq\mathrm{Cay}(\mathbb{F}_{q^{2}}^{+},E_{Q_{-}})$. The isomorphism is induced by the linear transformation providing the equivalence between $Q$ and $Q_{\pm}$. It is well-known that both $G_{+}$ and $G_{-}$ are (almost) Ramanujan graphs. This follows from the character tables of affine association schemes and Deligne’s bound for Kloosterman sums (e.g., [1, 14, 6]).
We summarize the connection to the Mesnager-Qi-Tang codes as follows:
Proposition 18.
For any odd prime power $q$, we have $E_{Q_{\pm}}=H_{\pm}$, and hence, the codes $\mathcal{C}(\mathbb{F}_{q}^{2},H_{\pm})$ in Theorem 14 coincide with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{\pm}})$. Furthermore, for any non-degenerate quadratic form $Q$, the code $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with either $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{+}})$ or $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{-}})$ up to coordinate permutations and sign changes (of coordinates), depending on whether $Q$ is equivalent to $Q_{+}$ or $Q_{-}$ in $\mathbb{F}_{q}$.
This proposition, combined with Theorem 14, implies that for any non-degenerate quadratic form $Q$, the set $E_{Q}$ essentially yields only the Mesnager-Tang-Qi codes. The second claim of Proposition 18 follows immediately from the Lemma below.
Lemma 19.
Let $Q$ and $Q^{\prime}$ be two non-degenerate quadratic forms over $\mathbb{F}_{q}$. If $Q$ and $Q^{\prime}$ are equivalent, then $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q^{\prime}})$ up to coordinate permutations and sign changes.
Proof.
By assumption, there exists $A\in\mathrm{GL}(2,\mathbb{F}_{q})$ such that $Q^{\prime}(\mathbf{x})=Q(A\mathbf{x})$ for all $\mathbf{x}\in\mathbb{F}_{q}^{2}$. This implies
$E_{Q^{\prime}}$ $=\{\mathbf{x}\in\mathbb{F}_{q}^{2}\mid Q^{\prime}(\mathbf{x})=1\}$
$=\{\mathbf{x}\in\mathbb{F}_{q}^{2}\mid Q(A\mathbf{x})=1\}$
$=\{A^{-1}\mathbf{y}\mid\mathbf{y}\in E_{Q}\}.$
Let $E_{Q}=S\sqcup(-S)$ where $S=\{\beta_{1},\dots,\beta_{n}\}$ is a fixed set of representatives. Then, the set $S^{\prime}=\{\beta^{\prime}_{1},\dots,\beta^{\prime}_{n}\}$ with $\beta^{\prime}_{i}=A^{-1}\beta_{i}$ forms a valid set of representatives for $E_{Q^{\prime}}$. Note that if $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q^{\prime}})$ is defined by another set of representatives $T^{\prime}$, then $S^{\prime}$ and $T^{\prime}$ differ only by the ordering of elements and the signs of coordinates (since $E_{Q^{\prime}}$ is inverse-closed).
Let $q=p^{k}$, where $p$ is an odd prime and $k\geq 1$. Viewing the ambient space $\mathbb{F}_{q}^{2}$ as a $2k$-dimensional vector space over $\mathbb{F}_{p}$, the map $A^{-1}$ acts as an $\mathbb{F}_{p}$-linear transformation, denoted by $f_{A}^{-1}$. A vector $\mathbf{c}=(c_{1},\dots,c_{n})\in\mathbb{F}_{p}^{n}$ is a codeword of $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ if and only if
$\sum_{i=1}^{n}c_{i}\beta_{i}=\mathbf{0}\quad(\text{in }\mathbb{F}_{p}^{2k}).$
Applying $f_{A}^{-1}$ to both sides yields
$f_{A}^{-1}\left(\sum_{i=1}^{n}c_{i}\beta_{i}\right)=\sum_{i=1}^{n}c_{i}(f_{A}^{-1}(\beta_{i}))=\sum_{i=1}^{n}c_{i}\beta^{\prime}_{i}=\mathbf{0}.$
This implies that $\mathbf{c}$ is a codeword for the code defined by representatives $\{\beta^{\prime}_{i}\}$. Since the representatives of $E_{Q^{\prime}}$ are unique only up to signs and ordering, the result follows. ∎
Example 20.
Let $Q(x,y)=x^{2}+y^{2}$.
  • If $q\equiv 1\pmod{4}$, then 1-1 is a square in $\mathbb{F}_{q}$, so let $i^{2}=-1$. $Q$ is equivalent to $Q_{-}$ via the transformation:
    $Q(x,y)=(x+iy)(x-iy)=Q_{-}(u,v),$
    where $u=x+iy$ and v=xiyv=x-iy. Thus, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{-}})$ up to coordinate permutations and sign changes.
  • If $q\equiv 3\pmod{4}$, then 1-1 is non-square. $Q$ is equivalent to $Q_{+}$ by setting $\delta=-1$:
    $Q(x,y)=x^{2}-(-1)y^{2}=Q_{+}(x,y).$
    Thus, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{+}})$.
Example 21.
Let $Q(x,y)=x^{2}+xy+y^{2}$. The discriminant of $Q$ is 3-3. Completing the square, we have
$Q(x,y)=\left(x+\frac{y}{2}\right)^{2}+\frac{3}{4}y^{2}.$
Let $u=x+\frac{y}{2}$ and $v=y$. Then $Q(x,y)$ is equivalent to the diagonal form $u^{2}+\frac{3}{4}v^{2}$.
  • If 3-3 is a square in $\mathbb{F}_{q}$, then $\frac{3}{4}$ can be written as $s^{2}$ for some $s\in\mathbb{F}_{q}^{*}$. Thus, $Q$ is equivalent to $u^{2}-s^{2}v^{2}=(u-sv)(u+sv)$, which is equivalent to $Q_{-}(x,y)=xy$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{-}})$ up to coordinate permutations and sign changes.
  • If 3-3 is a non-square in $\mathbb{F}_{q}$, then $\frac{3}{4}$ is a non-square (since $\frac{1}{4}$ is a square). Let $\delta=-\frac{3}{4}$. Then $Q$ is equivalent to $u^{2}-\delta v^{2}$, which is exactly $Q_{+}(u,v)$. In this case, $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q})$ coincides with $\mathcal{C}(\mathbb{F}_{q}^{2},E_{Q_{+}})$ up to coordinate permutations and sign changes.
6. Concluding remarks
  • It would be of significant interest to investigate whether other abelian (almost) Ramanujan graphs harbor generating sets that yield new families of quasi-perfect Lee codes. Such discoveries would further solidify the profound connection between the theory of Lee codes and spectral graph theory.
  • The 2-quasi-perfect Lee codes presented in this paper, as well as those in [15], are derived from abelian (almost) Ramanujan graphs that are spectrally indistinguishable from random graphs of the same edge density. That is, their normalized eigenvalue distributions converge to the semicircle distribution, a hallmark of typical random graphs. In this sense, these graphs exhibit strong pseudo-randomness. Conversely, the property of 2-quasi-perfectness imposes rigid structural regularities on the underlying Cayley graph, as dictated by Lemma 2. Notably, the strictly deterministic condition in Lemma 2 (2) is not a property that random graphs possess with high probability. Therefore, we observe that spectral indistinguishability does not preclude the existence of the highly ordered local structures required for $2$-quasi-perfectness. This underscores a fascinating dichotomy: spectral indistinguishability alone is insufficient to guarantee “local” pseudo-randomness, revealing a structural gap akin to the phenomena discussed in [7].
Acknowledgment
The author deeply appreciates anonymous referees for their valuable comments. This work was supported by JSPS KAKENHI Grant Number JP 23K13007. This work was also supported by Institute of Mathematics for Industry, Joint Usage/Research Center in Kyushu University (FY2025 Short-term Joint Research, Reference No. 2025a037).